OpenStudy (anonymous):

integrate t sec^2 (2t) dt

OpenStudy (anonymous):

integrate sec^2(2t) dt, integrate by parts, set dg=sec^2(2t)dt and g=(1/2)tan(2t), then you get (1/2) t tan(2t)-(1/2) integral (tan(2t) dt). Do you manage the rest?

OpenStudy (anonymous):

\[\int\limits_{}^{} t*\sec^{2}(2t)dt = t*\tan (2t) - \int\limits_{}^{}\tan (2t)dt \]

OpenStudy (anonymous):

this is going to be: = \[t*\tan (2t) + \ln |\cos (t)| + c\]

OpenStudy (anonymous):

all of this was done by parts

OpenStudy (anonymous):

all good except for typo in last line. it of course should be \[\ln(\cos(2t))\]

OpenStudy (anonymous):

in fact we are missing some constants too. should have \[\frac{1}{2}t \tan(2t)+\frac{1}{4} \ln(\cos(2t))\] if i am not mistaken

OpenStudy (anonymous):

@dumcow did i get this right? second constant i think is 1/4 because you need the 1/2 from first part when you do parts yes?

OpenStudy (anonymous):

the constants always throw me off so i could easily be wrong

OpenStudy (anonymous):

Satellite, I would agree with you over just about anyone on this site :) So I would probably agree your answer is right xP

OpenStudy (anonymous):

And you're right if I'm doing this in my head correctly lol.